Graphically show the variation of (i) ‘g’ with r for r>Re, r
(ii) Gravity with r for r>Re r

Grade:11

1 Answers

Vikas TU
5 years ago
Dear student,
For r greater than Re,
g = GM/(R + r)^2 => GM/R^2(1 + r/R)^2 = gs/(1+h/R)^2 (where gs = GM/R^2)
Further, g = gs*(1+h/R)^-2 = > gs*(1 – 2h/R)
 
For r less than Re, it would be
g  = GM(R – r)/R^3 => gs*(1-h/R)